Hi,Please may I ask for help?
What is the rate constant of a first-order reaction that takes 234 seconds for the reactant concentration to drop to half of its initial value?
1
Expert's answer
2020-06-04T10:38:50-0400

The time taken for the reactant concentration to drop to half of its initial value is called the half-life of the reaction "t_{1\/2}". For the first order reaction (A"\rightarrow" products), the rate law is:

"[A] = [A]_0e^{-kt}" .

At half-life:

"\frac{[A]}{[A]_0} = 1/2" .

Therefore, the rate constant of this reaction is:

"0.5 = e^{-kt_{1\/2}}"

"\text{ln}(0.5) = -kt_{1/2}"

"\text{ln}(2) = kt_{1/2}"

"k = \frac{\text{ln}(2)}{t_{1/2}} = \frac{\text{ln}(2)}{234} = 2.96·10^{-3}" s-1.

Answer: The rate constant of this first-order reaction is 2.96·10-3 s-1.